You have the following information on an economy.
C = 0.8Yd
I = GHS78
G = GHS70
TR = GHS40
T = 0.25Y
Derive an expression for aggregate expenditure based on this model.
Determine the ste
1
Expert's answer
2021-09-08T09:30:52-0400

"AE = C + I + G + TR"


"C = 0.8 (Y-T)"

"C = 0.8Y - 0.8(0.25)Y"

"C = 0.8 Y -0.2 Y" "= 0.6Y"

"AE = 70 + 40 + 78 + 0.6Y"

"AE = 188 + 0.6Y"